Maths

A speaks truth in 70% cases and B in 85% cases. In Wa

percen tage of cases are they likely to contradict each other in stating

the same fact.
Answer #1

P(A)=P [A Speaks truth] =70/100

P(B)=p(B speaks truth )=85/100

P(A')=p( A does not speak truth )=30/100

P(B')=p(B does not speak truth)=15/100

=P(they contradict each other )

=P(A speaks truth and B lies or B speak truth and A lies ) 

= P( A and B' or B and A')

=P(A)P(B)+P(B)(A')

=(70/100×15/100)+85/100×30/100)=1050/10000+2550/10000=.36

Percentage of cases in which they contradict =.36×100=36%

answered by: Vijayan